OpenStudy (anonymous):

A student bought 11 books for class. If the cost of his anatomy book was 3 times the mean cost of the other 10 books, then the cost of the anatomy book was what fraction of the total amount he paid for the 11 books? Can anyone show me how to do this algebraically instead of plugging in numbers?

OpenStudy (kropot72):

Let the mean cost of the other 10 books = m Cost of anatomy book = 3m Total amount paid for the 11 books = m + 3m = 4m Can you finish it now?

OpenStudy (anonymous):

I apologize i forgot the answer choices. Maby this should make more sense: A) 2/13 B) 3/13 C) 3/11 D) 3/10 E) 10/13

OpenStudy (anonymous):

If I'm not wrong when I do it the way you explained i get 3/4

OpenStudy (kropot72):

Sorry for my mistake. The total amount paid for the 11 books = 3m + 10m = 13m Can you find the answer now?

OpenStudy (anonymous):

why is that 10m?

OpenStudy (kropot72):

The average cost of each of the other 10 books = m Therefore the total cost of the other 10 books is the average cost of each of the other 10 books multiplied by the number of books = m * 10 = 10m Note that the anatomy book (the eleventh one) is not included in this part of the cost calculation.

OpenStudy (kropot72):

The total amount paid for the 11 books = 3m + 10m = 13m The amount paid for the anatomy book is 3m

OpenStudy (anonymous):

got it thanx :)

OpenStudy (kropot72):

You're welcome :)